ashauniwilliams06:

Write an equation in slope-intercept form for the line described. Write the slope and y-intercept as fractions, if necessary. passes through (5, –3), perpendicular to y = –2x + 4

ashauniwilliams06:

please help me with this

Vocaloid:

It’s perpendicular to y = -2x + 4. That means the function you’ll be writing has the *opposite reciprocal* slope as y = -2x + 4. First identify the slope of y = -2x + 4 (Recall y = mx + b form where m is the slope) Then take the opposite reciprocal (-1/m) This is the slope for the new function. It passes through (5, -3) so you can set up another y = mx + b equation, this time letting m be the new slope and (x,y) be the point (5, -3). Solve for b. Finally write the equation in y = mx + b form, using the new slope and b value.